every evening jenna empties out her pockets and puts her change in a jar. at the end of the week she counts her money. one week she had 40 coins, all of them dimes and quaters. when she added them up she had a total of &7.75.
how many dimes did jenna have??

Answer by Earlsdon(6294)

Let Q = the number of quarters and D = the number of dimes.
Then:
1) Q+D = 40 "One week she had 40 coins, all of them dimes and quarters...."
2) 0.25Q+0.1D = $7.75 "...she had a total of $7.75."
Rewrite equation 1) as: D = 40-Q and substitute into equation 2).
2a) 0.25Q+0.1(40-Q) = 7.75 Simplify the left side.
2b) 0.25Q+4-0.1Q = $7.75 Subtract 4 from both sides.
2c) 0.25Q-0.1Q = 3.75 Combine the like-terms on the left side.
2d) 0.15Q = 7.75 Divide both sides by 0.15
2e) Q = 25 and:
D = 40-Q
D = 40-25
D = 15
Jenna had 25 quarters and 15 dimes.